A mother's age is three less than 3 times her daughters age. The sum of there ages is 45. Find their ages
Answer by rimjhim.tomar(21)

let the age of the daughter be x yrs.
then the mother's age would be= 3x-3
given= sum of their ages= 45
x+3x-3 =45
4x-3 =45
4x =48
x =12
so the daughter's age= x= 12 yrs
and mother's age = 3x-3 = 36-3 = 33yrs